# Kinetic theory of dilute granular gases with hard-core and inverse power-law potentials under uniform shear flow: comparison with simplified model

### v1

@@ Abstract
We develop a kinetic-theory framework to investigate the steady rheology of a dilute gas interacting via a repulsive potential under uniform shear flow. Starting from the Boltzmann equation with a restitution coefficient that depends on the impact velocity and potential strength, we derive evolution equations for the stress tensor based on Grad’s moment expansion. The resulting expressions for the collisional rates and transport coefficients are fitted with simple analytical functions that capture their temperature dependence over a wide range of shear rates. Comparison with direct simulation Monte Carlo (DSMC) results shows excellent quantitative agreement for the shear stress, temperature anisotropy, and steady viscosity. We also analyze the velocity distribution functions, revealing that the system remains nearly Maxwellian even under strong shear.
@@ IV Boltzmann equation
where $\mathcal{E}$ is the macroscopic restitution coefficient given in Eq. (7).
@@
$\left(\frac{\partial}{\partial t}-\dot{\gamma}\frac{\partial}{\partial V_{1,x}}\right)f(\bm{V}_{1},t)=J(f,f),$ (12)
which corresponds to the standard form of the Boltzmann equation in the uniform-shear frame. Here, the collision integral implicitly contains information about the repulsive potential through $\mathcal{E}$ and $\theta_{\alpha}$. In the dilute limit considered, the stress tensor $P_{\alpha\beta}$ is determined purely by kinetic contributions, allowing direct comparison with kinetic-theory results for uncharged systems.
Multiplying both sides of Eq. (9) by $mV_{1,\alpha}V_{1,\beta}$ and integrating over $\bm{V}_{1}$, we obtain [25, 26, 27, 28, 18, 20]
$\frac{\partial}{\partial t}P_{\alpha\beta}+\dot{\gamma}\left(\delta_{\alpha x}P_{y\beta}+\delta_{\beta x}P_{\alpha y}\right)=-\Lambda_{\alpha\beta},$ (13)
@@
$\Lambda_{\alpha\beta}\equiv-\int\mathrm{d}\bm{V}_{1}mV_{\alpha}V_{\beta}J(f,f),$ (14)
@@
$P_{\alpha\beta}\equiv\int\mathrm{d}\bm{V}_{1}mV_{\alpha}V_{\beta}f(\bm{V},t),$ (15)
is the $\alpha\beta$ component of the stress tensor. It should be noted that, since we focus only on dilute systems, there is no contribution to the stress from particle contacts. Using Einstein’s summation convention $P_{\alpha\alpha}=\sum_{\alpha=1}^{3}P_{\alpha\alpha}$ and denoting the number density by $n$, the hydrostatic pressure satisfies $P=P_{\alpha\alpha}/3=nT$, where the temperature $T$ is given by
@@
$\frac{\partial P_{xy}}{\partial t}$ $=-\dot{\gamma}\left(T-\frac{1}{3}\Delta T\right)-\Lambda_{xy},$ (17c)
@@
Once the set of Eqs. (17) is solved, the rheological properties of the system can be determined. However, the explicit form of $\Lambda_{\alpha\beta}$ remains unknown at present. For this reason, we adopt Grad’s approximation for the velocity distribution function [29, 30, 31, 25, 26, 27, 24, 28, 18, 20]
$f(\bm{V},t)\approx f_{\mathrm{M}}(\bm{V},t)\left[1+\frac{m}{2T}\left(\frac{P_{\alpha\beta}}{nT}-\delta_{\alpha\beta}\right)V_{\alpha}V_{\beta}\right],$ (18)
@@
which is known to be a good approximation at least for hard-sphere and harmonic potential systems. Under this assumption, $\Lambda_{\alpha\beta}$ can be expressed explicitly as [25, 26, 27, 28, 18, 20]
$\Lambda_{\alpha\beta}=\zeta nT\delta_{\alpha\beta}+\nu\left(P_{\alpha\beta}-nT\delta_{\alpha\beta}\right),$ (20)
@@
The temperature dependence of $\Omega_{\alpha,n}^{(1)}$ and $\Omega_{\alpha,n}^{(2)}$ is shown in Fig. 5. Interestingly, $\Omega_{\alpha,n}^{(i)}$ exhibits only a weak dependence on the choice of the restitution coefficient $e$. This can be interpreted as follows: at high temperatures, almost all collisions are inelastic, whereas at low temperatures they are almost entirely elastic. In the intermediate regime, a gradual transition occurs. Since both limiting behaviors coincide, the dependence on $e$ does not become pronounced even in this crossover region.
@@
Based on these results, in the following we focus on the case of restitution coefficient $e=0.9$ and examine the corresponding rheological behavior.
@@ V Steady rheology
Figure 6 shows the dependence of the temperature, temperature anisotropy, and viscosity on the shear rate.
@@
We also present the direct simulation Monte Carlo (DSMC) [32, 33] results in Fig. 6. (For details of the DSMC method, see, e.g., Ref. [34], among others.) These results show good agreement over a wide range of shear rates, demonstrating the broad applicability of our theoretical framework.
@@ V.1 Introduction of a fitting model
Here, we propose a simple fitting form for $\Omega_{\alpha,n}^{(i)}$ as
@@
$\Omega^{(2)\mathrm{fit}}_{\alpha,7}$ $=\Omega^{(2)\mathrm{HC}}_{\infty,7}\left\{1+\log\left[1+b_{3}\left(\frac{T}{\varepsilon}\right)^{-a_{3}}\right]\right\},$ (25c)
where $a_{i}$ and $b_{i}$ ($i=1,2,3$) are fitting parameters. By applying the least-squares method, these coefficients are determined as
$a_{1}$ $=0.773,\ b_{1}=0.232,\ a_{2}=0.817,$
$b_{2}$ $=-0.107,\ a_{3}=2.04,\ b_{3}=0.134.$ (26)
Figure 7 compares the original data of $\Omega^{(i)\mathrm{fit}}_{\infty,n}$ with the proposed fitting functions. Although slight deviations are observed in $\Omega^{(1)\mathrm{fit}}_{\infty,n}$ ($i=1,2$) around the crossover region between the low- and high-temperature regimes, the overall agreement is satisfactory.

### v2

# Kinetic theory of dilute weakly charged granular gases with hard-core and inverse power-law interactions under uniform shear flow

@@ Abstract
We develop a kinetic-theory framework to investigate the steady rheology of a dilute gas interacting via a repulsive potential under uniform shear flow. Starting from the Boltzmann equation with a restitution coefficient that depends on the impact velocity and potential strength, we derive evolution equations for the stress tensor based on Grad’s moment expansion. The resulting expressions for the collisional rates and transport coefficients are fitted with simple analytical functions that capture their temperature dependence over a wide range of shear rates. Comparison with direct simulation Monte Carlo (DSMC) results shows excellent quantitative agreement for the shear stress, temperature anisotropy, and shear viscosity. We also analyze the velocity distribution functions, revealing that the system remains nearly Maxwellian even under strong shear.
@@ I Introduction
In the present work, we focus on the regime of weakly charged granular particles, where mechanical inelastic contact remains the dominant dissipation mechanism in binary collisions, while the electrostatic interaction acts as a perturbative long-range correction. In other words, the typical electrostatic potential energy at contact is assumed to be smaller than or comparable to the granular kinetic energy. Strongly charged regimes, in which long-range Coulomb interactions suppress direct mechanical contact or lead to collective plasma-like effects, are beyond the scope of the present kinetic description.
@@ IV Boltzmann equation
where $\mathcal{E}$ is the macroscopic restitution coefficient given in Eq. (7). Here, we do not describe the detailed microscopic dynamics during contact. Instead, we assume a scattering picture in which the asymptotic pre- and post-collisional states at infinite separation are connected by an effective collision rule characterized by the macroscopic restitution coefficient $\mathcal{E}$. Under this assumption, the mapping between the pre-collisional velocities $\{\bm{v}_{1}^{\prime\prime},\bm{v}_{2}^{\prime\prime}\}$ and the post-collisional velocities $\{\bm{v}_{1},\bm{v}_{2}\}$ given by Eq. (11) is one-to-one and linear in the normal component of the relative velocity. Therefore, the Jacobian of the transformation $\{\bm{v}_{1}^{\prime\prime},\bm{v}_{2}^{\prime\prime}\}\to\{\bm{v}_{1},\bm{v}_{2}\}$ can be explicitly evaluated, yielding the factor $1/\mathcal{E}^{2}$ in the gain term. As a result, the collision operator is formally equivalent to that of inelastic hard spheres, with the constant restitution coefficient $e$ replaced by the effective coefficient $\mathcal{E}$.
@@
$\left(\frac{\partial}{\partial t}-\dot{\gamma}V_{1,y}\frac{\partial}{\partial V_{1,x}}\right)f(\bm{V}_{1},t)=J(f,f),$ (12)
which corresponds to the standard form of the Boltzmann equation in the uniform-shear frame. Here, the collision integral implicitly contains information about the repulsive potential through $\mathcal{E}$ and $\theta_{\alpha}$. In the dilute limit considered, the stress tensor $P_{k\ell}$ is determined purely by kinetic contributions, allowing direct comparison with kinetic-theory results for uncharged systems.
Multiplying both sides of Eq. (9) by $mV_{1,k}V_{1,\ell}$ and integrating over $\bm{V}_{1}$, we obtain [25, 26, 27, 28, 18, 20]
$\frac{\partial}{\partial t}P_{k\ell}+\dot{\gamma}\left(\delta_{kx}P_{y\ell}+\delta_{\ell x}P_{ky}\right)=-\Lambda_{k\ell},$ (13)
@@
$\Lambda_{k\ell}\equiv-\int\mathrm{d}\bm{V}_{1}mV_{1,k}V_{1,\ell}J(f,f),$ (14)
@@
$P_{k\ell}\equiv\int\mathrm{d}\bm{V}mV_{k}V_{\ell}f(\bm{V},t),$ (15)
is the $k\ell$ component of the stress tensor. It should be noted that, since we focus only on dilute systems, there is no contribution to the stress from particle contacts. Using Einstein’s summation convention $P_{k}=\sum_{k=x,y,z}^{3}P_{kk}$ and denoting the number density by $n$, the hydrostatic pressure satisfies $P=P_{kk}/3=nT$, where the temperature $T$ is given by
@@
$\frac{\partial P_{xy}}{\partial t}$ $=-\dot{\gamma}n\left(T-\frac{1}{3}\Delta T\right)-\Lambda_{xy},$ (17c)
@@
Once the set of Eqs. (17) is solved, the rheological properties of the system can be determined. However, the explicit form of $\Lambda_{k\ell}$ remains unknown at present. For this reason, we adopt Grad’s approximation for the velocity distribution function [29, 30, 31, 25, 26, 27, 24, 28, 18, 20]
$f(\bm{V},t)\approx f_{\mathrm{M}}(\bm{V},t)\left[1+\frac{m}{2T}\left(\frac{P_{k\ell}}{nT}-\delta_{k\ell}\right)V_{k}V_{\ell}\right],$ (18)
@@
which is known to be a good approximation at least for hard-sphere and harmonic potential systems. Under this assumption, $\Lambda_{k\ell}$ can be expressed explicitly as [25, 26, 27, 28, 18, 20]
$\Lambda_{k\ell}=\zeta nT\delta_{k\ell}+\nu\left(P_{k\ell}-nT\delta_{k\ell}\right),$ (20)
@@
The temperature dependence of $\Omega_{\alpha,n}^{(1)}$ and $\Omega_{\alpha,n}^{(2)}$ is shown in Fig. 5. Interestingly, $\Omega_{\alpha,n}^{(i)}$ exhibits only a weak dependence on the choice of the restitution coefficient $e$. This can be interpreted as follows: at high temperatures, almost all collisions are inelastic, whereas at low temperatures they are almost entirely elastic. In the intermediate regime, a gradual transition occurs. Since both limiting behaviors coincide, the dependence on $e$ does not become pronounced even in this crossover region. Based on these results, in the following we focus on the case of restitution coefficient $e=0.9$ and examine the corresponding rheological behavior.
@@ V Steady rheology
Before proceeding, it is important to clarify the physical condition under which the steady uniform shear flow (USF) is realized. The steady state is determined from the energy balance condition, namely, that the viscous heating term $2/(3n)\dot{\gamma}P_{xy}$ is exactly balanced by the collisional contribution $\Lambda_{\alpha\alpha}/(3n)$. Unlike the conventional inelastic hard-sphere model with constant restitution coefficient, the collisional moment $\Lambda_{\alpha\alpha}$ in the present model is not a simple function of temperature, since it reflects the velocity dependence of the effective restitution coefficient. Therefore, for given values of the core restitution coefficient (at $r=d$) and the imposed shear rate $\dot{\gamma}$, the steady temperature $T$ is determined implicitly from the energy balance equation. In this sense, $\dot{\gamma}$ and the restitution parameter are externally controlled and can be regarded as independent, while the reduced shear rate $\dot{\gamma}^{*}$ follows from the resulting steady-state temperature.
Figure 6 shows the dependence of the temperature, temperature anisotropy, and viscosity on the shear rate. Here, we have employed the dimensionless quantities as
$T^{*}\equiv\frac{T}{\varepsilon},\ \Delta T^{*}\equiv\frac{\Delta T}{\varepsilon},\ \dot{\gamma}^{*}\equiv\dot{\gamma}d\sqrt{\frac{m}{\varepsilon}},\ \eta^{*}\equiv\eta\frac{d^{2}}{\sqrt{m\varepsilon}}.$ (24)
@@
We also present the direct simulation Monte Carlo (DSMC) [32, 33] results in Fig. 6. (For details of the DSMC method, see, e.g., Ref. [34].) These results show good agreement over a wide range of shear rates, demonstrating the broad applicability of our theoretical framework.
Figure 7: Shear rate dependence of the the viscosity for $\alpha=4$, $6$, 88, 1010, 1212 when we fix $e=0.9$. The inset shows a magnified view of the low-shear region.
Figure 7 shows the shear rate dependence of the reduced viscosity for fixed $e=0.9$ and several values of the inverse-power-law exponent $\alpha=4$, $6$, 88, 1010, and $12$. As in Fig. 6, the results become almost independent of $\alpha$ in the regime $\dot{\gamma}^{*}\gtrsim 0.2$, corresponding to $T\gtrsim\varepsilon$. In this high-shear regime, the kinetic energy dominates over the repulsive potential barrier, and the system approaches the hard-core limit. Consequently, the influence of the exponent $\alpha$ becomes negligible and the flow curves asymptotically approach the Bagnold scaling. In contrast, a clear dependence on $\alpha$ emerges in the low-shear regime $\dot{\gamma}^{*}\lesssim 0.2$. In this region, the viscosity increases with increasing $\alpha$. This behavior reflects the stronger suppression of low-velocity collisions as the inverse-power-law interaction becomes steeper, which enhances the effective resistance to shear.
@@ V.1 Introduction of a fitting model
The evaluation of Eq. (22) requires a double integration for each temperature. Therefore, when Eq. (22) must be evaluated repeatedly, for instance in iterative calculations of transport coefficients, the computational cost becomes significant.
To overcome this difficulty, we introduce simple fitting forms for $\Omega_{\alpha,n}^{(i)}$ that preserve the correct asymptotic behaviors in both the low- and high-temperature limits:
@@
$\Omega^{(2)\mathrm{fit}}_{\alpha,7}$ $=\Omega^{(2)\mathrm{HC}}_{\infty,7}\left\{1+\log\left[1+b_{3}\left(\frac{T}{\varepsilon}\right)^{-a_{3}}\right]\right\}.$ (26c)
The hyperbolic tangent form for $\Omega^{(1)}_{\alpha,n}$ is chosen so as to reproduce the vanishing behavior in the low-temperature regime and the hard-sphere limit $\Omega^{(1)\mathrm{HC}}_{\infty,n}$ in the high-temperature regime, while ensuring a smooth crossover between the two limits. Similarly, the functional form adopted for $\Omega^{(2)}_{\alpha,7}$ correctly interpolates between the temperature-dependent behavior at low temperatures and the constant asymptotic value at high temperatures.
The fitting parameters $a_{i}$ and $b_{i}$ ($i=1,2,3$) are determined by the least-squares method, yielding
$a_{1}$ $=0.773,\quad b_{1}=0.232,\quad a_{2}=0.817,$
$b_{2}$ $=-0.107,\quad a_{3}=2.04,\quad b_{3}=0.134.$ (27)
The fitted curves reproduce the numerical integration results with high accuracy over the entire temperature range considered as shown in Fig. 8.